# Designing reaction-cross-diffusion systems with Turing and wave instabilities

### v1

@@ 5 Discussion
We have developed an approach to designing reaction-cross-diffusion systems exhibiting Turing or wave instabilities for systems involving several components. We showed how to constructively build unstable principal submatrices which essentially encode these instabilities, and place these within the design constraints of a given system. The assumptions in Theorems 1 and 2 are rather mild, suggesting that this process can be employed for large classes of systems, leaving flexibility in the constructions to allow for additional design constraints if they arise. This demonstrates some of the freedom gained in allowing non-diagonal diffusion tensors, or in having complete control over linearized kinetics. Finally, we showed these ideas in four distinct models with a range of linear and nonlinear transport, highlighting further issues of model construction that may arise, such as in preventing blowups or ensuring the positivity of solutions when this is required. We remark that all but the first example demonstrated employing these ideas in cases where we did not allow complete freedom over either matrix in the linearization, but instead had to work in a more constrained setting.
@@
Although the linear theory developed is useful, Turing and wave instabilities do not ensure on their own that one will always be able to find specific kinds of solutions emerging from the instability of a spatially homogeneous equilibrium. We focused on the instability of a single homogeneous steady state, but realistic systems may admit more than one such state, leading to cases where the linear theory fails to capture the emergence of patterning due to other attracting states Andrewturing . As we showed in the first example, calculations of the criticality of the bifurcations can be invaluable to ensure that a nearby stable pattern exists in the case of a supercritical bifurcation. In contrast, a subcritical bifurcation is a situation where the linear theory is not always a good indicator of subsequent emergent behaviour FahadWoods ; villar2023degenerate ; burke2007homoclinic ; Villar-sepulveda-wave . We note, in particular, that the two-species Keller-Segel model with logistic growth can exhibit spatiotemporal chaos after a subcritical Turing instability, despite such two-species systems not exhibiting wave instabilities or even Hopf bifurcations painter2011spatio . Hence, there is substantial work to be done in building design principles which extend to the weakly nonlinear setting, and even developing a language of constraints for nonlinearities.
In a different direction, we note that, depending on the purpose of the mathematical model, having too much freedom in the design can lead to unsatisfying theories. Given full control over the nonlinearities, one can design even two-component reaction-diffusion systems that exhibit a wide variety of pre-defined behaviours woolley2021bespoke , with essentially the freedom to match any kind of patterned state. Given such freedom, one must impose constraints that involve both linear and nonlinear aspects of the system under study to ensure that the choices made make sense in the modelling context. Similarly, while the results we have presented are general and powerful, one can argue that they provide too much freedom. In particular, there are few systems where one has complete control over either transport or reactions. Nevertheless, experimental protocols are increasingly providing more ways to create and tune a number of systems to exhibit patterning horvath2009experimental ; karig2018stochastic ; konow2021insights , and the fundamental routes to instabilities provided here can provide some guidance on pursuing these in increasingly sophisticated systems even under constraints. Importantly, we note that the ideas used apply even with more constraints in the given system, as shown in the examples.
@@ Appendix A Useful results for designing unstable systems
Let $\mathbf{\hat{D}}$ be a diagonalizable matrix with non-negative real eigenvalues. Assume that $\mathbf{\hat{D}}$ has $\ell$ eigenvalues $\sigma_{i_{1}},\ldots,\sigma_{i_{\ell}}$ that tend to $\infty$, while all the others remain positive but finite. Then, there exist $\ell$ eigenvalues of $\mathbf{J}-\mu\mathbf{\hat{D}}$ having real parts that tend to $\infty$- ∞ for every $\mu>0$, while $n-\ell$ eigenvalues of said matrix tend to the eigenvalues of the principal submatrix of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ formed out of the rows and columns that are complementary to $i_{1},\ldots,i_{\ell}$, as $\mu\to 0^{+}$.
@@
which has eigenvalues $-9.50278\pm 81.3221i,-0.29445$, meaning it is stable. We note that it has a $2\times 2$ unstable principal submatrix at the bottom-right corner with eigenvalues $0.0513167$ and $1.94868$, with an element within this submatrix being 3. In particular, if we consider a diffusion matrix given by
@@
the dispersion relation of a potential system we can generate with these matrices for different values of $\delta$ that range from 0 to 0.002 in uniform steps of size 0.0004 is shown in Figure 14. Here, we note that there is a strong interaction between Turing and wave instabilities. In fact, when $\delta=0$, the largest eigenvalue converges to a real number as $k\to\infty$, as it is expected. However, for larger values of $\delta$, there is an interaction between real and complex-conjugate eigenvalues as there is a range of values of said parameter under which two complex conjugate eigenvalues split into two real eigenvalues for a range of values of $k$. This implies that several different scenarios can be generated through different unstable submatrices in the Jacobian matrix of the system. The question is then how can we generate a stable matrix from an unstable one? Part of the answer is provided in the following lemma, which once again, allows for several degrees of freedom.
Figure 14: Dispersion relation of a system whose linearization is given by (34) and the diffusion matrix is set as (35).
@@
with $\mathbf{v}_{m}=\left(v_{m,1},\ldots,v_{m,\ell+1}\right)^{\intercal}$ for each $11\leq m\leq\ell+$ 1, then there exists a set of real numbers $\left\{c_{1,s}\right\}_{s=1}^{\ell+1}$ such that the matrix
@@
Proof.
As the eigenvalues of $\mathbf{A}$ are all different, then a matrix is similar to $\mathbf{A}$ if and only if it has the same characteristic polynomial as $\mathbf{A}$. By Lemma 8, the characteristic polynomial of $B$ is given by
@@
$b_{0}=\det\left(\mathbf{B}\right)$, and $b_{\ell+1}=(-1)^{\ell+1}$. Observe that the first two sums in (37) are linear combinations of the coefficients $\left\{c_{1,s}\right\}_{s=1}^{\ell+1}$ whilst the third one is a constant term depending only on $\mathbf{K}$.
@@
Note that, by assumption, there exist coefficients $\left\{c_{s,1}\right\}_{s=2}^{\ell+1}$ so that the matrix of coefficients of this system of equations is invertible, which implies the existence of a solution $\left\{c_{1,s}\right\}_{s=1}^{\ell+1}$ so that the conclusion holds. ∎

### v2

@@ 5 Discussion
We have developed an approach to designing reaction-cross-diffusion systems exhibiting Turing or wave instabilities for systems involving several components. We showed how to constructively build unstable principal submatrices which essentially encode these instabilities, and place these within the design constraints of a given system. The assumptions in Theorems 1 and 2 are rather mild, suggesting that this process can be employed for large classes of systems, leaving flexibility in the constructions to allow for additional design constraints if they arise. This demonstrates some of the freedom gained in allowing non-diagonal diffusion tensors, or in having complete control over linearized kinetics. Finally, we showed these ideas in four distinct models with a range of linear and nonlinear transport, highlighting further issues of model construction that may arise, such as in preventing blowups or ensuring the positivity of solutions when this is required. We remark that all but the first example demonstrated employing these ideas in cases where we did not allow complete freedom over either matrix in the linearization, but instead had to work in a more constrained setting. An important thermodynamic consideration is to restrict to diffusion matrices $\mathbf{\hat{D}}$ that are symmetric, as these can more easily be justified from the thermodynamic perspective klika2024onsager . In such cases, Theorem 2 can be applied directly, although the constructions used in proving Theorem 1 would require significant changes that we leave for future work.
@@
Although the linear theory developed is useful, Turing and wave instabilities do not ensure on their own that one will always be able to find specific kinds of solutions emerging from the instability of a spatially homogeneous equilibrium. We focused on the instability of a single homogeneous steady state, but realistic systems may admit more than one such state, leading to cases where the linear theory fails to capture the emergence of patterning due to other attracting states Andrewturing . As we showed in the first example, calculations of the criticality of the bifurcations can be invaluable to ensure that a nearby stable pattern exists in the case of a supercritical bifurcation. In contrast, a subcritical bifurcation is a situation where the linear theory is not always a good indicator of subsequent emergent behaviour FahadWoods ; villar2023degenerate ; burke2007homoclinic ; Villar-sepulveda-wave . We note, in particular, that the two-species Keller-Segel model with logistic growth can exhibit spatiotemporal chaos after a subcritical Turing instability, despite such two-species systems not exhibiting wave instabilities or even Hopf bifurcations of homogeneous states painter2011spatio (though see ei2014spatio ; kong2024existence for cases where patterned states undergo Hopf bifurcations leading to these spatiotemporal behaviours). Hence, there is substantial work to be done in building design principles that extend to the weakly nonlinear setting, and even developing a language of constraints for nonlinearities. Related to this limitation, the assumption of homogeneous Neumann or periodic conditions precludes the interplay of more complex boundary conditions which can modify patterns and even initiate patterning outside of the Turing space determined here via linear theory maini1997boundary ; krause2021isolating .
In a different direction, we note that, depending on the purpose of the mathematical model, having too much freedom in the design can lead to unsatisfying theories. Given full control over the nonlinearities, one can design even two-component reaction-diffusion systems that exhibit a wide variety of pre-defined behaviours woolley2021bespoke ; woolley2025bespoke , with essentially the freedom to match any kind of patterned state. Given such freedom, one must impose constraints that involve both linear and nonlinear aspects of the system under study to ensure that the choices made make sense in the modelling context. Similarly, while the results we have presented are general and powerful, one can argue that they provide too much freedom. In particular, there are few systems where one has complete control over either transport or reactions. Nevertheless, experimental protocols are increasingly providing more ways to create and tune a number of systems to exhibit patterning horvath2009experimental ; karig2018stochastic ; konow2021insights , and the fundamental routes to instabilities provided here can provide some guidance on pursuing these in increasingly sophisticated systems even under constraints. Importantly, we note that the ideas used apply even with more constraints in the given system, as shown in the examples.
@@ Appendix A Useful results for designing unstable systems
Let $\mathbf{\hat{D}}$ be a diagonalizable matrix with non-negative real eigenvalues. Assume that $\mathbf{\hat{D}}$ has $\ell$ eigenvalues $\sigma_{i_{1}},\ldots,\sigma_{i_{\ell}}$ that tend to $\infty$, while all the others remain positive but finite. Then, there exist $\ell$ eigenvalues of $\mathbf{J}-\mu\mathbf{\hat{D}}$ having real parts that tend to -$\infty$ for every $\mu>0$, while $n$-$\ell$ eigenvalues of said matrix tend to the eigenvalues of the principal submatrix of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ formed out of the rows and columns that are complementary to $i_{1},\ldots,i_{\ell}$, as $\mu\to 0^{+}$.
@@
which has eigenvalues $9.50278\pm 81.3221i,-0.29445$, meaning it is stable. We note that it has a $2\times 2$ unstable principal submatrix at the bottom-right corner with eigenvalues $0.0513167$ and $1.94868$, with an element within this submatrix being 3. In particular, if we consider a diffusion matrix given by
@@
the dispersion relation of a potential system we can generate with these matrices for different values of $\delta$ that range from 0 to 0.002 in uniform steps of size 0.0004 is shown in Fig. 14. Here, we note that there is a strong interaction between Turing and wave instabilities. In fact, when $\delta=0$, the largest eigenvalue converges to a real number as $k\to\infty$, as it is expected. However, for larger values of $\delta$, there is an interaction between real and complex-conjugate eigenvalues as there is a range of values of said parameter under which two complex conjugate eigenvalues split into two real eigenvalues for a range of values of $k$. This implies that several different scenarios can be generated through different unstable submatrices in the Jacobian matrix of the system. The question is then how can we generate a stable matrix from an unstable one? Part of the answer is provided in the following lemma, which once again, allows for several degrees of freedom.
Figure 14: Dispersion relation of a system whose linearization is given by Eq. 34 and the diffusion matrix is set as Eq. 35.
@@
with $\mathbf{v}_{m}=\left(v_{m,1},\ldots,v_{m,\ell+1}\right)^{\intercal}$ for each $1\leq m\leq\ell+1$, then there exists a set of real numbers $\left\{c_{1,s}\right\}_{s=1}^{\ell+1}$ such that the matrix
@@
Proof: As the eigenvalues of $\mathbf{A}$ are all different, then a matrix is similar to $\mathbf{A}$ if and only if it has the same characteristic polynomial as $\mathbf{A}$. By Lemma 8, the characteristic polynomial of $B$ is given by
@@
$b_{0}=\det\left(\mathbf{B}\right)$, and $b_{\ell+1}=(-1)^{\ell+1}$. Observe that the first two sums in Eq. 37 are linear combinations of the coefficients $\left\{c_{1,s}\right\}_{s=1}^{\ell+1}$ whilst the third one is a constant term depending only on $\mathbf{K}$.
@@
Note that, by assumption, there exist coefficients $\left\{c_{s,1}\right\}_{s=2}^{\ell+1}$ so that the matrix of coefficients of this system of equations is invertible, which implies the existence of a solution $\left\{c_{1,s}\right\}_{s=1}^{\ell+1}$ so that the conclusion holds.